# Spectroscopy of two-dimensional interacting lattice electrons using symmetry-aware neural backflow transformations

### v1

@@ II.1 Fermions on the lattice
It will prove useful to connect the two main formalisms for describing fermionic systems: first quantization, which labels the fermions, and second quantization where we consider the occupation number basis or a given orbital set (here the lattice sites). To establish the correspondence, we consider a canonical ordering of the lattice sites through their label assignment $i=\{1,...,N_{s}\}$. The latter can be chosen arbitrarily, and in practice we choose a snake-like ordering in the case of the 2D lattice. This enables us to recover the particle positions in the first quantization framework from a given occupation number configuration in second quantization (see Ref. [22]). We introduce $x=(\mathbf{r}_{i_{1}},\mathbf{r}_{i_{2}},\ldots,\mathbf{r}_{i_{N_{f}}})$, where $i_{p}$ is the site index occupied by the $p^{\text{th}}$ electron, where the fermion number index is determined by the chosen canonical ordering. In other words, we have an ordered set of indices $i_{1}<i_{2}<..<i_{N_{f}}$. Furthermore, we introduce the occupation number configuration $n=(n_{1},\ldots,n_{N_{s}})\in\{0,1\}^{N_{s}}$. Hence, in this notation, the canonical ordering allows us to extract active or occupied lattice indices $\{i_{p}\}_{p=1,\ldots,N_{f}}$, given the occupation numbers $n$. This establishes implicit mappings $x=x(n)$ and $n=n(x)$ (see Fig. 1).
Figure 1: Implicit mappings of $x(n)$ and $n(x)$. This figure shows how the occupation configuration $n=(n_{1},\ldots,n_{N_{s}})=(0,0,1,1,0,0,0,1,0)$ maps to the position vectors $x=(\mathbf{r}_{i_{1}},\mathbf{r}_{i_{2}},\ldots,\mathbf{r}_{i_{N_{f}}})$, and vice versa, illustrating the relationship between occupied lattice sites and their indices respecting a canonical ordering.
II.2 Wavefunction ansatz
II.2.1 Slater-Jastrow
@@
The two-body Jastrow factor is defined as [29, 47]
@@
II.2.2 Neural Backflow Transformation
The Slater-Jastrow wavefunction ansatz can be further improved by including backflow transformations, thereby significantly increasing the expressiveness of the model. We use the neural backflow transformation that effectively promotes the single-particle orbitals to many-body orbitals [39, 26, 27]. Therefore, we introduce the backflow function $F$ that produces a configuration-dependent orbital matrix $F\in\mathbb{C}^{N_{f}\times N_{s}}$. We then adapt the orbital matrix as
@@ II.2.2 Neural Backflow Transformation
II.3 Symmetries and Excitations
We consider an electronic Hamiltonian on a lattice that commutes with the elements of a symmetry group $G$, such as total spin, the total momentum, and geometrical symmetries such as rotation. The eigenstates of these the many-body Hamiltonian can be classified with the symmetry sectors of $G$. We restrict the NQS ansatz to a given symmetry sector labeled by $I$ through a quantum-number projection
@@ II.3 Symmetries and Excitations
where $\mathbf{K}$ is the total momentum and the sum runs over all possible translation vectors. We use the above-mentioned quantum number projection to compute both the ground state and the low-lying excited states. The low-lying excitations are characterized by different momentum sectors, and their computation involves optimizing the wavefunction within the quantum number sectors distinct from the ground state [43, 45]. Instead of focusing on a single excited state individually, an alternative strategy involves adopting a multi-target approach, which has recently been introduced for continuous systems [48]. However, for cost-effectiveness and interpretation in terms of translation quantum numbers, we focus on the symmetry-projection method outlined above.
In a brute-force approach, the evaluation of the symmetrized wave function $\psi^{I}(n)$ for a configuration $n$ would require $\left|G\right|$ evaluations of the parametrized non-symmetric wave function $\psi$ in Eq. (6). In particular, for translation symmetry in Eq. (7) of a square lattice of size $L\times L$, this would require $\left|G\right|=L^{2}$ evaluations. The computational burden induced by the symmetrization procedure can therefore become significant for increasing system sizes. For this purpose, we introduce a novel set of symmetry-aware neural backflow transformations that require only a single evaluation of a neural network to produce all $\psi(\hat{g}^{-1}n)$ (i.e. $\forall g\in G$) required to evaluate the projection in Eq. (6). This will allow us to reach larger system sizes, even when considering deep neural networks to represent the backflow transformation. In the next section, we discuss the requirements of this symmetry-aware backflow transformation and introduce a specific neural parametrization to fulfill the constraints.
II.3.1 Equivariance Condition
@@
Our key objective is to preserve translation equivariance in the backflow transformation. Using this, we can construct a symmetrized Neural Slater-Backflow-Jastrow ansatz, given that the Jastrow correlation function is translation-invariant and the backflow is constructed as a neural network respecting the equivariance conditions in Eqs. (8) and (9). A natural candidate for a symmetry-equivariant neural network is a convolutional neural network (CNN) operating on occupation configurations [49, 50, 51], which naturally exhibits these properties. In a CNN, spatial translations in the input lead to corresponding shifts in the output feature maps. The occupation configurations undergo multiple CNN-transformation layers, resulting in an output of the same size $L^{2}$ as the input configuration. We employ $N_{f}$ independent backflow transformations corresponding to the different orbitals $\mu$. From the outputs, we obtain the reduced matrix $\overline{B}$ in Eq. (5), by selecting the columns corresponding to the occupied sites from the resulting backflow matrix $F_{\mu,i}(n)$. We depict this procedure in Fig. 2 where we provide a visual representation of a CNN backflow satisfying the equivariance conditions. In summary, given our CNN backflow is equivariant: instead of evaluating the CNN for all elements of the translation symmetry group, we can extract all the required $\psi\left(\hat{g}^{-1}n\right)$ in Eq. (6) from the output of a single evaluation of the backflow CNN. This approach improves efficiency and reduces computational redundancy in handling symmetrical transformations. In Fig. 2 we explicitly show this symmetry-averaging process under equivariance conditions. For additional information on the architecture and its adaptation to different system sizes, we refer to Appendix A.
@@
Figure 2: Visual representations of the symmetry-aware backflow, the translational equivariance condition, and the quantum number projection. Panel a) showcases how the backflow determinants are constructed. The neural network $\mathrm{CNN}_{\mu}$ for a given orbital $\mu$ takes as input the configuration $n$, and produces a backflow vector output $F_{\mu}(n)$. The reduced matrix $\bar{F}_{\mu}$ is obtained by selecting the active indices $i$ from $F_{\mu}$. These active indices are linked to the occupied sites, represented by the dark green blocks in the input $n$, using the canonical ordering. Panel b) demonstrates the equivariance property of the backflow function. Applying a symmetry transformation to the input as $\hat{g}^{-1}n$ and then extracting the active indices, we get $\bar{F}_{\mu}(\hat{g}^{-1}n)$. This is equivalent to applying the symmetry transformation directly to the active indices represented by $\hat{g}i$. This results equivalently in the reduced matrix $\bar{F}_{{\mu\hat{g}}}(n)$. Panel c) represents the quantum number projection. Following a single CNN evaluation, we obtain the reduced matrices $\bar{F}_{\mu g}$, $\forall\hat{g}\in G$ by properly constructing the reduced matrix, without reevaluating the backflow CNN. Subsequently, we compute the symmetry-averaged wavefunction $\psi(\hat{g}^{-1}n)$.
III Results
III.1 Hamiltonian and Observables
The Hamiltonian of the $t$-$V$ model reads
The first term describes electron hopping between neighboring sites with hopping parameter $t$. The second term corresponds to the nearest neighbor Coulomb repulsion with interaction strength $V\geq 0$. We will set $t=1$ from hereon. We further decompose the Hilbert space $\mathcal{H}$ into fixed particle-number subspaces $\mathcal{H}_{N_{f}}$ [52]. The $t$-$V$ model was originally introduced to study the thermodynamic and transport properties of superconductors [53, 54]. Additionally, it provides a conceptual framework for explaining phenomena such as phase separation or stripe order in cuprates and organic conductors [55, 56, 57, 58, 59, 60]. In practice, the $t$-$V$ model can be realized, for example, in experiments with strongly polarized ${}^{3}\mathrm{He}$ atoms  [61, 62]. Despite its apparent simplicity, the $t$-$V$ model cannot be solved analytically in two or higher dimensions. It also reveals highly nontrivial phase transitions that have been studied in previous works with various computational techniques, including variational Monte Carlo [63, 64, 55, 65, 66, 67, 68, 22, 69, 70].
In the strong-coupling limit where $V/t\rightarrow\infty$ we encounter a charge-ordering (CO) phase where the system behaves classical. At large $V/t$, it is energetically unfavorable for two fermions to occupy neighboring lattice sites. The correlations become short-ranged, suggesting localization and insulating behavior, giving rise to a charge-ordered insulating phase. At half-filling ($\bar{n}=0.5$) the charge order corresponds to a checkerboard pattern. Conversely, at weak interaction strengths $V/t$, the fermions can easily hop between neighboring lattice sites, and the system behaves like a non-interacting free Fermi gas. In this weak coupling limit the system enters the metallic phase and becomes exactly solvable at $V/t\rightarrow 0$ [63, 64, 71, 22, 72, 73, 74, 75, 76].
We introduce the normalized density-density correlation function as [22]
where $S_{R}(i)=\{j\in\mathcal{V}:d(i,j)=R\}$ is the set of vertices with a fixed distance $R$ from the vertex $i$. Another important observable to detect the CO phase is the structure factor [63, 64, 66, 67, 77]:
In the CO phase, well-defined peaks at $\mathbf{K}=(\pi,\pi)$ indicate a checkerboard charge pattern, and in the thermodynamic limit $S((\pi,\pi))/N_{s}$ converges to a finite value, reflecting long-range order. We will study the $t$-$V$ model on a two-dimensional square lattice of side $L$ and with periodic boundary conditions, for various system sizes $|\mathcal{V}|=L^{2}$ and different interaction strengths $V/t$, with densities close to half-filling and at closed momentum shells.
III.2 Ground States
We benchmark our symmetrized Neural Slater-Backflow-Jastrow ($\psi^{K}_{BF}$) ansatz with respect to a mean-field Slater determinant (which is equivalent to Hartree-Fock (HF)), a symmetrized Slater-Jastrow ($\psi^{K}$) without backflow, and a non-symmetrized Slater-Jastrow. Furthermore, we compare with ground state energies obtained with another state-of-the-art neural quantum state method (“Slater-Jastrow with an additional sign correction neural network”) from Ref. [22]. In Fig. 3 (a), we show results for a small system size ($L=4$), allowing comparison to results obtained by exact diagonalization (ED). Mean-field Slater relative errors range from $10^{-3}$ to $10^{-1}$, with accuracies decreasing at large interaction strengths $V/t$. Our symmetrized backflow ansatz consistently yields ground-state errors below $10^{-4}$, also at higher values of $V/t$. We also observe that the symmetry-aware backflow transformation yields the most accurate ground-state energies over the whole interaction regime. For a larger system size ($L=10$), shown in Fig. 3 (b), this trend is confirmed. The backflow corrections significantly lower the estimated ground-state energy across all coupling strengths. We compare the converged VMC energies with HF energies by computing the difference $E-E_{\text{HF}}$. Our results show that the backflow ansatz consistently provides lower energies across the full range of interaction strengths.
Figure 3: Comparison of ground-state energies of the $t$-$V$ model obtained with various wavefunction models. We evaluate the performance of our symmetrized Neural Slater-Backflow-Jastrow ($\psi^{K}_{BF}$) in comparison to a symmetrized Slater-Jastrow ($\psi^{K}$) without backflow, a non-symmetrized Slater-Jastrow, and the neural quantum state “Slater-Jastrow with an additional sign correction neural network” described in Ref. [22]. In panel a), we present results for $L=4$ with $\bar{n}=0.44$ and show relative errors compared to exact diagonalization (ED) as a function of the interaction strengths $V/t$. In panel b), we show the deviation from Hartree-Fock (HF) energies for a $L=10$ system with filling fraction $\bar{n}=0.44$ defined as $E-E_{HF}$.
In Fig. 4 (a), we show the density-density correlation function as defined in Eq. 11 for $L=10$ close to half-filling $\bar{n}=0.44$ (see Appendix B for results on the $8\times 8$ system). When the interaction strength increases, the correlations start to oscillate more pronouncedly as a result of the increasingly ordered charge distribution. In the CO phase, the amplitudes of the oscillations decrease more gradually with distance compared to that of the metallic phase. For weak couplings, the correlations barely oscillate and fade as the graph distance $R$ increases.
To pinpoint the transition point in the thermodynamic limit, we use finite-size scaling. We study the structure factor $S(\pi,\pi)/N_{s}$ for various $V/t$. The critical transition point is found where the structure factor attains a finite value in the thermodynamic limit. In Fig. 4 (b) we plot the structure factor $S(\pi,\pi)/N_{s}$ as a function of the size of the system $1/L$ (for $L=4,6,8,10$) and in Table 1 we report the extrapolated results in the thermodynamic limit. Near half-filling, specifically for $\bar{n}=0.44$, we estimate that the transition occurs at $V_{c}/t\simeq 1.14\pm 0.04$, which is consistent with the value reported in Ref. [22].
Figure 4: a) The two-point density-density correlation functions as a function of graph distance $R$ for different $V/t$ for a system size of $L=10$ and $\bar{n}=0.44$. b) Finite-size scaling of the structure factor $S(\pi,\pi)/N_{s}$ versus different system sizes $1/L$ (where $L=4,6,8,10$) with $\bar{n}=0.44$ for different values of $V/t$ close to the estimated critical point $V_{c}/t\simeq 1.14\pm 0.04$.
$V/t$ $S_{\infty}$
0.1 $-0.026\pm 0.001$
0.6 $-0.021\pm 0.002$
1.0 $-0.003\pm 0.001$
1.3 $0.003\pm 0.001$
1.6 $0.042\pm 0.002$
2.7 $0.085\pm 0.002$
Table 1: Structure factor extrapolations to the thermodynamic limit (taken from Figure 4 (b)) as a function of $V/t$.
III.3 Excitations
[…]
Figure 5: Single quadrant of the first Brillouin zone of the two-dimensional square lattice with high-symmetry points $(k_{u},k_{v})$.
To capture low-lying excitations, we carry out VMC optimizations across various momentum sectors $\mathbf{K}=(k_{u},k_{v})$, where $k_{u,v}=\{0,\pm\frac{2\pi m}{L_{u,v}}\}$, and $m=1,..,\frac{L_{u,v}}{2}$. Here $L_{u}$ and $L_{v}$ are the side lengths of the two-dimensional lattice in the $\mathbf{e}_{u}$ and $\mathbf{e}_{v}$ directions. In Fig. 5, we represent a single quadrant within the corresponding first Brillouin zone with conventional symbols to represent high-symmetry points  [78].
Figure 6: (top panel) The lowest and second-lowest energies for the designated $\mathbf{K}$ sector using ED, compared to the lowest energies in each sector, obtained with our variational ansatzes. (bottom panel) The relative energy errors associated with the ansatzes. These results are for a closed-shell system with L = 4 and a density of $\bar{n}=0.31$.
We first benchmark the performance of our symmetrized Neural Slater-Backflow-Jastrow variational ansatz ($\psi^{K}_{BF}$) and a symmetrized Slater-Jastrow ($\psi^{K}$) (top) model for a system size of $L=4$ and $\bar{n}=0.31$ in Fig. 6. We observe that also in symmetry sectors different from the ground state, the symmetric backflow transformation reduces the relative energy error. We show the lowest energy state in each sector and compare it with results obtained from ED. In the lower panel of Fig. 6, we calculate the corresponding relative errors with respect to the ED energies. For our backflowed ansatz ($\psi^{K}_{BF}$) these lie between $10^{-6}-10^{-3}$ for all sectors. In Appendix D, we also include additional simulation data of the open shell $L=4$ system (see Fig. 13).
Prior studies have documented a range of coexistence phenomena for large and finite $V/t$ away from half-filling, transitioning from phase separation to potential stripe and checkerboard coexistence. In Ref. [79], the $t$-$V$ model on a square lattice with nearest-neighbor repulsive interactions was studied using mean-field theory for small system sizes. The authors observed a second-order phase transition from the Fermi liquid to the $(\pi,\pi)$ charge density wave state. At stronger repulsion, charge density waves coexisted at different momentum sectors when doped away from half-filling. In Ref. [58] and the subsequent study in Ref. [59], exact diagonalization was used to study small 2D systems. They found that at high repulsion and around quarter-filling densities, doped holes formed stable charged stripes acting as anti-phase walls [80], which are stable against phase separation in fermionic systems.
We extend our analysis to larger system sizes $L=8$ with $\bar{n}=0.39$ and $L=10$ with $\bar{n}=0.41$ to reduce finite-size effects. We simulate each system size with distinct particle density to ensure the absence of ground-state degeneracies in the non-interacting limit. We confirm the persistent non-degenerate ground state at the $\Gamma=(0,0)$ point (see Fig. 7 for $L=8$ and Fig. 12 in Appendix C for $L=10$). By comparing our symmetrized Neural Slater-Backflow-Jastrow ($\psi^{K}_{BF}$) ansatz with a symmetrized Slater-Jastrow ($\psi^{K}$) without backflow, we observe improved energy levels, even for low-lying excited states, with the inclusion of the backflow correction term.
Figure 7: Lowest excitation energies in different $\mathbf{K}$ sectors for an $L=8$ closed-shell system with $\bar{n}=0.39$ across a range of $V/t$ values.
Figure 8: Energy gaps $\Delta K$ between the ground state (located at the $\Gamma$ point) and excitation energies $(K=M,X)$ for a) $L=8$ lattice with filling factor $\bar{n}=0.39$ and b) $L=10$ lattice with filling factor $\bar{n}=0.41$, both being closed-shell systems with no ground state degeneracy. We include $\Delta K_{\infty}$ for $V/t=\infty$ in both figures, showing a collapse to a fully localized, charge-ordered phase at infinite repulsion.
III.4 V-score
We now aim to generalize the assessment of the performance of our model. Since ED becomes intractable for larger system sizes, we rely on the recently introduced V-score as a guiding metric [81], which can be computed using the variational energy and its variance. The V-score is dimensionless and invariant under energy shifts by construction. It is defined as [81]:
The V-score serves as a valuable tool for discerning which Hamiltonians and regimes pose challenges for arbitrary classical variational techniques, even when we lack prior knowledge about the precise exact solution. Its practicality lies in its ability to quantify the accuracy of a particular method independently, without the need for direct comparisons with other methods. In particular, this metric enables us to draw comparisons between the accuracy obtained with our method on the given Hamiltonian, compared to other commonly studied condensed matter Hamiltonians (including spin Hamiltonians).
Figure 9: Ground state V-scores comparison for different ansatz: $\psi^{K}_{BF}$ (backflow), $\psi^{K}$ (Slater-Jastrow), and HF (Hartree-Fock), across system sizes $L=4$ with $\bar{n}=0.31$, and $L=10$ with $\bar{n}=0.41$.
Figure 10: V-score for a symmetrized Neural Slater-Backflow-Jastrow as a function of $V/t$ for different ground (GS) and excited states for system sizes $L=4$ with $\bar{n}=0.31$, $L=8$ with $\bar{n}=0.39$, and $L=10$ with $\bar{n}=0.41$.
IV Conclusion and Outlook
In this work, we introduced a novel approach to studying the low-energy excitation spectrum of fermionic Hamiltonians. By introducing symmetry-aware neural backflow transformations, we show that we can target the eigenstates of fermionic Hamiltonians with high accuracy. As a benchmark comparison, we show that this approach also yields significantly more accurate ground-state energies than other state-of-the-art variational Monte Carlo approaches.
In particular, we introduce equivariance conditions for the backflow transformations that lead to an efficient symmetry projection. We show that convolutional neural networks yield a powerful parametrization for our symmetry-aware backflow transformations that fulfill thee equivariance conditions for both translation and particle-permutation symmetry. This key contribution enables us to efficiently access excited states by varying the total momentum $\mathbf{K}$ in the quantum number projection. Furthermore, we have showcased the utility of our approach in identifying phase-transitions on the $t$-$V$ model at system sizes far beyond what is reachable with exact diagonalization. To this end, we computed correlation functions and structure factors, resulting in the pinpointing of the critical point at $V_{c}/t=1.14$. We also computed the V-score to quantify the variational accuracy of our proposed ansatz for different interaction regimes, system sizes and excitations. Previous analysis based on the V-score has highlighted the challenging nature of targeting fermionic eigenstates. Our observations indicate that the symmetry-aware backflow ansatz yields accurate ground states and performs favorably compared to other state-of-the-art methods over the full interaction regime, including when strong correlations occur. Additionally, we find that our method yields accurate approximations to the low-energy eigenstates with a given $K$ momentum.
Future extensions of our approach include generalizing it by including additional symmetries, such as rotational and reflection symmetries. This will involve using more general group-convolutional kernels, as in group-convolutional neural networks (GCNN) [49, 50, 51]. We focused on the nearest neighbor $t$-$V$ model, but an extension is to consider Hamiltonians where spin-degrees of freedom become relevant as well (such as the Fermi-Hubbard model), or where interactions beyond nearest neighbor terms become relevant.

### v2

@@ IV.1 Fermions on the lattice
It will prove useful to connect the two main formalisms for describing fermionic systems: first quantization, which labels the fermions, and second quantization where we consider the occupation number basis or a given orbital set (here the lattice sites). To establish the correspondence, we consider a canonical ordering of the lattice sites through their label assignment $i=\{1,...,N_{s}\}$. The latter can be chosen arbitrarily, and in practice we choose a zigzag-like ordering in the case of the 2D lattice. This enables us to recover the particle positions in the first quantization framework from a given occupation number configuration in second quantization (see Ref. Stokes_2020). We introduce $x=(\mathbf{r}_{i_{1}},\mathbf{r}_{i_{2}},\ldots,\mathbf{r}_{i_{N_{f}}})$, where $i_{p}$ is the site index occupied by the $p^{\text{th}}$ electron, where the fermion number index is determined by the chosen canonical ordering. In other words, we have an ordered set of indices $i_{1}<i_{2}<..<i_{N_{f}}$. Furthermore, we introduce the occupation number configuration $n=(n_{1},\ldots,n_{N_{s}})\in\{0,1\}^{N_{s}}$. Hence, in this notation, the canonical ordering allows us to extract active or occupied lattice indices $\{i_{p}\}_{p=1,\ldots,N_{f}}$, given the occupation numbers $n$. This establishes implicit mappings $x=x(n)$ and $n=n(x)$ (see Fig. 9).
Figure 9: Implicit mappings between occupation $n(x)$ and position configurations $x(n)$. This figure shows how the occupation configuration $n=(n_{1},\ldots,n_{N_{s}})=(0,0,1,1,0,0,0,1,0)$ maps to the position vectors $x=(\mathbf{r}_{i_{1}},\mathbf{r}_{i_{2}},\ldots,\mathbf{r}_{i_{N_{f}}})$, and vice versa, illustrating the relationship between occupied lattice sites and their indices respecting a canonical ordering.
IV.2 Wavefunction ansatz
IV.2.1 Slater-Jastrow
@@
The two-body Jastrow factor is defined as Jastrow_1955, Pei_2021
@@
IV.2.2 Neural Backflow Transformation
The Slater-Jastrow wavefunction ansatz can be further improved by including backflow transformations, thereby significantly increasing the expressiveness of the model. We use the neural backflow transformation that effectively promotes the single-particle orbitals to many-body orbitals Luo_2019, Pfau_2020, Hermann_2020. Therefore, we introduce the backflow function $F$ that produces a configuration-dependent orbital matrix $F\in\mathbb{C}^{N_{f}\times N_{s}}$. We then adapt the orbital matrix as
@@ IV.2.2 Neural Backflow Transformation
IV.3 Symmetries and Excitations
We consider an electronic Hamiltonian on a lattice that commutes with the elements of a symmetry group $G$, such as total spin, the total momentum, and geometrical symmetries such as rotation. The eigenstates of the many-body Hamiltonian can be classified with the symmetry sectors of $G$. We restrict the NQS ansatz to a given symmetry sector labeled by $I$ through a quantum-number projection
@@ IV.3 Symmetries and Excitations
where $\mathbf{K}$ is the total momentum and the sum runs over all possible translation vectors. Due to the antisymmetric nature of the wavefunction, translating fermions can lead to a change in the sign of the wavefunction, depending on how their positions are permuted. This effect is naturally handled by the above-mentioned quantum number projection with corresponding characters, ensuring that the fully symmetrized wavefunction remains translationally invariant.
We use this approach to compute both the ground state and the low-lying excited states. The low-lying excitations are characterized by different momentum sectors, and their computation involves optimizing the wavefunction within the quantum number sectors distinct from the ground state Mizusaka_Imada_2004, Nomura_2021. Instead of focusing on a single excited state individually, an alternative strategy involves adopting a multi-target approach, which has recently been introduced for continuous systems Pfau_2023. However, for cost-effectiveness and interpretation in terms of translation quantum numbers, we focus on the symmetry-projection method outlined above.
In a brute-force approach, the evaluation of the symmetrized wave function $\psi^{I}(n)$ for a configuration $n$ would require $\left|G\right|$ evaluations of the parametrized non-symmetric wave function $\psi$ in Eq. (11). In particular, for translation symmetry in Eq. (12) of a square lattice of size $L\times L$, this would require $\left|G\right|=L^{2}$ evaluations. The computational burden induced by the symmetrization procedure can therefore become significant for increasing system sizes. For this purpose, we introduce a set of symmetry-aware neural backflow transformations that are constructed as equivariant functions, requiring only a single evaluation of a neural network to produce all $\psi(\hat{g}^{-1}n)$ (i.e. $\forall g\in G$) required to evaluate the projection in Eq. (11). This will allow us to reach larger system sizes, even when considering deep neural networks to represent the backflow transformation. In the next section, we discuss the requirements of this symmetry-aware backflow transformation and introduce a specific neural parametrization to fulfill the constraints.
IV.3.1 Equivariance Condition
@@
Our key objective is to preserve translation equivariance in the backflow transformation. Using this, we can construct a symmetrized Neural Slater-Backflow-Jastrow ansatz, given that the Jastrow correlation function is translation-invariant and the backflow is constructed as a neural network respecting the equivariance conditions in Eqs. (13) and (14). A natural candidate for a translation-equivariant neural network is a convolutional neural network operating on occupation configurations Cohen_2016, Roth_2021, Roth_2023, which naturally exhibits these properties. In a CNN, spatial translations in the input lead to corresponding shifts in the output feature maps, effectively preserving spatial locality. The occupation configurations undergo multiple CNN-transformation layers, resulting in outputs of the same size $L^{2}$ as the input configuration. This preserves the structure needed to maintain equivariance, without averaging over symmetry group elements as done for invariance. We employ $N_{f}$ independent backflow transformations corresponding to the different orbitals $\mu$. From the outputs, we obtain the reduced matrix $\overline{B}$ in Eq. (10), by selecting the columns corresponding to the occupied sites from the resulting backflow matrix $F_{\mu,i}(n)$. We depict this procedure in Fig. 10 where we provide a concise visual representation of a CNN backflow satisfying the equivariance conditions. In summary, given our CNN backflow is equivariant: instead of evaluating the CNN for all elements of the translation symmetry group, we can extract all the required $\psi\left(\hat{g}^{-1}n\right)$ in Eq. (11) from the output of a single evaluation of the backflow CNN. This approach improves efficiency and reduces computational redundancy in handling symmetrical transformations. In Fig. 10, we illustrate this symmetry-averaging process under equivariance conditions. For additional information on the architecture and its adaptation to different system sizes, we refer to Supplementary Note I.
@@
Figure 10: Representation of symmetry-aware backflow and quantum number projection. Panel a) showcases how the backflow determinants are constructed. The neural network $\mathrm{CNN}_{\mu}$ for a given orbital $\mu$ takes as input the configuration $n$, and produces a backflow vector output $F_{\mu}(n)$. The reduced matrix $\bar{F}_{\mu}$ is obtained by selecting the active indices $i$ from $F_{\mu}$. These active indices are linked to the occupied sites, represented by the dark green blocks in the input $n$, using the canonical ordering. Panel b) demonstrates the equivariance property of the backflow function. Applying a symmetry transformation to the input as $\hat{g}^{-1}n$ and then extracting the active indices, we get $\bar{F}_{\mu}(\hat{g}^{-1}n)$. This is equivalent to applying the symmetry transformation directly to the active indices represented by $\hat{g}i$. This results equivalently in the reduced matrix $\bar{F}_{{\mu\hat{g}}}(n)$. Panel c) represents the quantum number projection. Following a single CNN evaluation, we obtain the reduced matrices $\bar{F}_{\mu g}$, $\forall\hat{g}\in G$ by properly constructing the reduced matrix, without reevaluating the backflow CNN. Subsequently, we compute the symmetry-averaged wavefunction $\psi(\hat{g}^{-1}n)$.